the width of a rectangle is one half its length. The perimeter of the rectangle is 24 in. what are the width and length of the rectangle show how to work it out
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

the width of a rectangle is one half its length. The perimeter of the rectangle is 24 in. what are the width and length of the rectangle ---------------------------------------
W = (1/2)L
24 = 2(W+L)
----------------
Rearrange and substitute to get:
L = 2W
24 = 2(W + 2W)
6W = 24
W = 4 inches.
L = 2W = 8 inches
======================
Cheers,
Stan H.

Answer by checkley77(12844)

W=L/2
2L+2W=24
2L+2(L/2)=24
2L+L=24
3L=24
L=24/3
L=8 IN. IS THE LENGTH.
W=8/2
W=4 IN. IS THE WIDTH.
PROOF:
2*8+2*4=24
16+8
24
24=24